OpenStudy (anonymous):

I'm a little confused with this question. I understand how to use Laplace in this on both sides but when I apply it to 2x and -3y do they both become functions of s? x'=2x-3y, x(0)=8 y(0)=3

OpenStudy (anonymous):

I ended up with s*X(t)-x(0) = L{2x} + L{-3y} would it be correct to say that the left side would just be \[\frac{ 2 }{ s^2 } - \frac{3}{s^2}\]

OpenStudy (anonymous):

I mean the right side